In 7 years Deac will be twice as old as Irina. Four
years ago he was three times as old as Irina was
then. How old are each of them now?

Answer by haileytucki(390)



A + 7 = 2(B+7)
A + 7 = 2B + 14
A - 2B = 7
and
A - 4 = 3(B-4)
A - 4 = 3B - 12
3B - A = 8

adding the two equations together:

[A - 2B] + (3B - A) = [7] + (8)
[A-2B]+(3B-A)=[7]+(8)
Remove the parentheses that are not needed from the expression.
A-2B+3B-A=[7]+(8)
Since A and -A are like terms, add -A to A to get 0.
0-2B+3B=[7]+(8)
Since -2B and 3B are like terms, subtract 3B from -2B to get B.
B=[7]+(8)
Remove the parentheses that are not needed from the expression.
B=7+8
Add 8 to 7 to get 15.
B=15
Now we have B=115, we substitute it into the original equation.

A-2(15)=7
Multiply -2 by each term inside the parentheses.
A-30=7
Since -30 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 30 to both sides.
A=30+7
Add 7 to 30 to get 37.
A=37

Deac is 37 and Irina is 15.